# An operator splitting analysis of Wasserstein–Fisher–Rao gradient flows

### v1

@@ Appendix C Preservation of log-concavity
$\widehat{c}_{2}$ $=b\tan\left(\tan^{-1}\left(\tfrac{c_{1}}{b}\right)-b\tau_{2}\right)$
@@
and $\nabla^{2}\mathcal{F}_{2}\succ\alpha_{2}I$ where
$\alpha_{2}=\left(1-\frac{1}{2}e^{-\tau_{2}}\right)\alpha_{\pi}+e^{-\tau_{2}}\widehat{c}_{2}.$
Proceeding inductively for $i=3,4,\dots$, in a similar way with required conditions on $\tau_{i}$ and $\delta_{i}$, we obtain
$\alpha_{i}$ $=\left(1-\tfrac{1}{2}e^{-\tau_{i}}\right)\alpha_{\pi}+e^{-\tau_{i}}\widehat{c}_{i}=c_{i}+\frac{\alpha_{\pi}}{2},\quad i=1,2,3,\dots$ (C.13)
@@
$\widehat{c}_{i}$ $=b\tan\left(\tan^{-1}\left(\tfrac{c_{i-1}}{b}\right)-b\tau_{i}\right)$ (C.14)
$c_{i}$ $=(1-e^{-\tau_{i}})\frac{\alpha_{\pi}}{2}+e^{-\tau_{i}}\widehat{c}_{i}$ (C.15)
$c_{0}$ $=\alpha_{d}+\frac{\delta_{0}}{2}\alpha_{\pi}.$ (C.16)
@@
Similarly to the limiting analysis in the proof of Lemma 4.1, we have that (C.14)-(C.15) corresponds to a splitting scheme of the ODE
$\dot{c}_{t}=-c_{t}^{2}-c_{t}-b^{2}+\frac{\alpha_{\pi}}{2},$ (C.17)
@@
whereby for the $i$th iteration, the ODE $\dot{c}_{t}=-c_{t}^{2}-b^{2}$, initialised at $c_{{i-1}}$ is solved over time $\tau_{i}$ yielding $\widehat{c}_{i}$ and then the ODE $\dot{c}_{t}=-c_{t}+\alpha_{\pi}$ initialised at $\widehat{c}_{i}$ is solved over time $\tau_{i}$ yielding $c_{i}$. Finally, due to (C.13), $\alpha_{t}=c_{t}+\frac{\alpha_{\pi}}{2}$, is the log-concavity constant of $\mu_{t}$, the solution of the exact WFR PDE (2.1) at time $t$.
@@
We now make precise the conditions on $\tau_{i}$ which must be satisfied in order for the limit (C.17) to hold as $\max_{i}\tau_{i}\rightarrow 0$. Recall that $\tau_{i}$ must be chosen such that $\widehat{c}_{i}>0$, which using (C.12), is possible if $0<\tau_{i}<\tau_{i}^{\ast}$ where
$\tau_{i}^{\ast}$ $=\frac{1}{b}\tan^{-1}\left(\frac{c_{i-1}}{b}\right),\quad i=1,2,3,\dots$ (C.18)
@@
Notice that there exists a $\tau_{1}^{\ast}>0$ whenever $\alpha_{d},\delta_{0}>0$, which implies $\widehat{c}_{1}>0$ and $c_{1}>\left(1-e^{-\tau_{1}}\right)\frac{\alpha_{\pi}}{2}$ and $\tau_{2}^{\ast}>\frac{1}{b}\tan^{-1}\left(\frac{(1-e^{-\tau_{1}})\alpha_{\pi}}{2b}\right)>0$. This then implies $\widehat{c}_{2}>0$ and once again $c_{2}>\left(1-e^{-\tau_{2}}\right)\frac{\alpha_{\pi}}{2}$ and $\tau_{3}^{\ast}>\frac{1}{b}\tan^{-1}\left(\frac{(1-e^{-\tau_{2}})\alpha_{\pi}}{2b}\right)>0$. By induction, $\tau_{i}^{\ast}>\frac{1}{b}\tan^{-1}\left(\frac{(1-e^{-\tau_{i-1}})\alpha_{\pi}}{2b}\right)>0$ and since $\tau_{i-1}<\tau_{i-1}^{\ast}$, we can instead consider the recursion
$\tau^{\ast}_{i}=\frac{1}{b}\tan^{-1}\left(\frac{(1-e^{-\tau_{i-1}^{\ast}})\alpha_{\pi}}{2b}\right),\enskip i=1,2,3,\dots$ (C.19)
@@
It is not difficult to see that $\tau_{i}^{\ast}>0$ for all $i=1,2,3,\dots$. Its limit, $\tau_{\infty}^{\ast}$ can be found from the fixed point equation $\tau^{\ast}_{\infty}=g(\tau_{\infty}^{\ast})$ where $g(\tau):=\frac{1}{b}\tan^{-1}\left(\frac{(1-e^{-\tau})\alpha_{\pi}}{2b}\right)$, for which $t_{\infty}^{\ast}=0$ is a valid (but not unique) solution. Notice also that $0<\frac{dg(\tau)}{d\tau}=\frac{\alpha_{\pi}}{2b^{2}}\frac{e^{-\tau}}{1+\left(\frac{(1-e^{-\tau})\alpha_{\pi}}{2b}\right)^{2}}<\frac{\alpha_{\pi}}{2b^{2}}$ for all $\tau>0$. Under the assumed condition (4.4), $\frac{dg(0)}{d\tau}=\frac{\alpha_{\pi}}{2b^{2}}>0$, $\tau=0$ is an unstable fixed point, so that $\tau_{\infty}^{\ast}\neq 0$ since $\tau_{1}^{\ast}\neq 0$. To show that there exists at least one other $0<\tau_{\infty}^{\ast}<\infty$, consider $h(\tau):=g(\tau)-\tau$. First note that $h(\tau)\rightarrow$-$\infty$ as $\tau\rightarrow\infty$ since $g(\tau)<\frac{1}{b}\tan^{-1}\left(\frac{\alpha_{\pi}}{2b}\right)$ for all $\tau\geq 0$. Additionally, $h(0)=0$ and $\frac{dh(0)}{d\tau}>0$ and $\frac{dh(\tau)}{d\tau}\rightarrow-1$ as $\tau\rightarrow\infty$ monotonically.
Then by continuity of $h$ and $g$, there must exist a $\tau>0$ such that $h(\tau)=0$, i.e. there exists a $\tau_{\infty}^{\ast}>0$. Therefore, we may construct a sequence of step sizes $\{\tau_{i}\}_{i=1,2,3,}$ with $\tau_{i}<\min(\tau_{1}^{\ast},\tau_{\infty}^{\ast}\neq 0)$ for all $i=1,2,3,\dots$ as a discretisation of the interval $[0,T]$ for all $T>0$, for which the recursion (C.14)-(C.15) converges to the continuous time ODE (C.17) as $\max_{i}\tau_{i}\rightarrow 0$.
As a further confirmation of the need for condition (4.4), consider the case $\frac{\alpha_{\pi}}{2b^{2}}<1$. The recursion (C.19) is a decreasing sequence with $\tau_{\infty}^{\ast}=0$ by a standard fixed point argument. Furthermore, the bound on $\frac{dg(\tau)}{d\tau}$ implies $g$ is Lipschitz and so $|\tau_{i+1}^{\ast}-\tau_{i}^{\ast}|=|g(\tau_{i}^{\ast})-g(\tau_{i-1}^{\ast})|\leq\frac{\alpha_{\pi}}{2b^{2}}|\tau_{i}^{\ast}-\tau_{i-1}^{\ast}|$. Then since $g(0)=0$, it holds that $\tau_{i}^{\ast}\leq\left(\frac{\alpha_{\pi}}{2b^{2}}\right)^{i}\tau_{0}^{\ast}$. Since $\sum_{i=1}^{\infty}\left(\frac{\alpha_{\pi}}{2b^{2}}\right)^{i}=\frac{1}{1-\frac{\alpha_{\pi}}{2b^{2}}}$ whenever $\frac{\alpha_{\pi}}{2b^{2}}<1$, it holds that $\sum_{i=1}^{\infty}\tau_{i}<\sum_{i=1}^{\infty}\tau_{i}^{\ast}\leq\frac{1}{1-\frac{\alpha_{\pi}}{2b^{2}}}$, so that a uniform in time control does not follow from the above sequence of thresholds $\{\tau_{i}^{\ast}\}$.
Finally, to obtain a uniform in time log-concavity constant, consider the fixed point of (C.17),
$c_{\infty}=\frac{-1\pm\sqrt{1+4(\frac{\alpha_{\pi}}{2}-b^{2})}}{2},$ (C.20)
which has one real positive solution under condition (4.4). Note also that $c_{t}$ approaches $c_{\infty}$ monotonically, as is true for scalar, autonomous Riccati ODEs. Its explicit solution is trivially found by solving the separable ODE, which we include here for completeness. Denote by $c_{\infty,1}$ and $c_{\infty,2}$, the positive and negative solutions of (C.20), respectively. Then,
$\int_{c_{0}}^{c_{T}}\frac{1}{(c_{t}-c_{\infty,1})(c_{t}-c_{\infty,2})}dc_{t}$ $=-\int_{0}^{T}dt$
$\frac{(c_{T}-c_{\infty,1})(c_{0}-c_{\infty,2})}{(c_{T}-c_{\infty,2})(c_{0}-c_{\infty,1})}$ $=e^{-T(c_{\infty,1}-c_{\infty,2})}$
$(c_{T}-c_{\infty,1})$ $=l_{0}e^{-T(c_{\infty,1}-c_{\infty,2})}(c_{T}-c_{\infty,2})$
$c_{T}$ $=\frac{c_{\infty,1}-c_{\infty,2}l_{0}e^{-T(c_{\infty,1}-c_{\infty,2})}}{1-l_{0}e^{-T(c_{\infty,1}-c_{\infty,2})}}$
where $l_{0}=\frac{(c_{0}-c_{\infty,1})}{(c_{0}-c_{\infty,2})}$. Finally, noting that $\alpha_{T}=c_{T}+\frac{\alpha_{\pi}}{2}$ yields the result.

### v2

@@ Appendix C Preservation of log-concavity
$\widehat{c}_{2}$ $=b\tan\left(\tan^{-1}\left(\tfrac{c_{1}}{b}\right)-2b\tau_{2}\right)$
@@
and $\nabla^{2}\mathcal{F}_{2}\succ\alpha_{2}I$ where $\alpha_{2}=\left(1-\frac{1}{2}e^{-\tau_{2}}\right)\alpha_{\pi}+e^{-\tau_{2}}\widehat{c}_{2}$. Proceeding inductively for $i=3,4,\dots$, in a similar way with required conditions on $\tau_{i}$ and $\delta_{i}$, we obtain
$\alpha_{i}$ $=\left(1-\tfrac{1}{2}e^{-\tau_{i}}\right)\alpha_{\pi}+e^{-\tau_{i}}\widehat{c}_{i}=c_{i}+\frac{\alpha_{\pi}}{2},\quad i=1,2,3,\dots$ (3.14)
@@
$\widehat{c}_{i}$ $=b\tan\left(\tan^{-1}\left(\tfrac{c_{i-1}}{b}\right)-2b\tau_{i}\right)$ (3.15)
$c_{i}$ $=(1-e^{-\tau_{i}})\frac{\alpha_{\pi}}{2}+e^{-\tau_{i}}\widehat{c}_{i}$ (3.16)
$c_{0}$ $=\alpha_{d}+\frac{\delta_{0}}{2}\alpha_{\pi}.$ (3.17)
@@
Similarly to the limiting analysis in the proof of Lemma 4.1, we have that (3.15)-(3.16) corresponds to a splitting scheme of the ODE
$\dot{c}_{t}=-2c_{t}^{2}-c_{t}-2b^{2}+\frac{\alpha_{\pi}}{2},$ (3.18)
@@
whereby for the $i$th iteration, the ODE $\dot{c}_{t}=-2c_{t}^{2}-2b^{2}$, initialised at $c_{{i-1}}$ is solved over time $\tau_{i}$ yielding $\widehat{c}_{i}$ and then the ODE $\dot{c}_{t}=-c_{t}+\frac{\alpha_{\pi}}{2}$ initialised at $\widehat{c}_{i}$ is solved over time $\tau_{i}$ yielding $c_{i}$. Finally, due to (3.14), $\alpha_{t}=c_{t}+\frac{\alpha_{\pi}}{2}$, is the log-concavity constant of $\mu_{t}$, the solution of the exact WFR PDE (2.1) at time $t$, when $c_{t}>0$.
@@
We now make precise the conditions on $\tau_{i}$ such that (3.18) has a strictly positive solution for any $t>0$. Recall that $\tau_{i}$ must be chosen such that $\widehat{c}_{i}>0$, which using (3.13), is possible if $0<\tau_{i}<\tau_{i}^{\ast}$ where
$\tau_{i}^{\ast}$ $=\frac{1}{2b}\tan^{-1}\left(\frac{c_{i-1}}{b}\right),\quad i=1,2,3,\dots$ (3.19)
@@
Notice that there exists a $\tau_{1}^{\ast}>0$ whenever $\alpha_{d},\delta_{0}>0$, which implies $\widehat{c}_{1}>0$ and $c_{1}>\left(1-e^{-\tau_{1}}\right)\frac{\alpha_{\pi}}{2}$ and $\tau_{2}^{\ast}>\frac{1}{2b}\tan^{-1}\left(\frac{(1-e^{-\tau_{1}})\alpha_{\pi}}{2b}\right)>0$. This then implies $\widehat{c}_{2}>0$ and once again $c_{2}>\left(1-e^{-\tau_{2}}\right)\frac{\alpha_{\pi}}{2}$ and $\tau_{3}^{\ast}>\frac{1}{2b}\tan^{-1}\left(\frac{(1-e^{-\tau_{2}})\alpha_{\pi}}{2b}\right)>0$. By induction, $\tau_{i}^{\ast}>\frac{1}{2b}\tan^{-1}\left(\frac{(1-e^{-\tau_{i-1}})\alpha_{\pi}}{2b}\right)>0$ and since $\tau_{i-1}<\tau_{i-1}^{\ast}$, we can instead consider the recursion
$\tau^{\ast}_{i}=\frac{1}{2b}\tan^{-1}\left(\frac{(1-e^{-\tau_{i-1}^{\ast}})\alpha_{\pi}}{2b}\right),\enskip i=1,2,3,\dots$ (3.20)
@@
It is not difficult to see that $\tau_{i}^{\ast}>0$ for all $i=1,2,3,\dots$. Its limit, $\tau_{\infty}^{\ast}$ can be found from the fixed point equation $\tau^{\ast}_{\infty}=g(\tau_{\infty}^{\ast})$ where $g(\tau):=\frac{1}{2b}\tan^{-1}\left(\frac{(1-e^{-\tau})\alpha_{\pi}}{2b}\right)$, for which $\tau_{\infty}^{\ast}=0$ is a valid (but not unique) solution. Now $g(\tau)$ is a Lipschitz function since $0<\frac{dg(\tau)}{d\tau}=\frac{\alpha_{\pi}}{4b^{2}}\frac{e^{-\tau}}{1+\left(\frac{(1-e^{-\tau})\alpha_{\pi}}{2b}\right)^{2}}\leq\frac{\alpha_{\pi}}{4b^{2}}$ for all $\tau\geq 0$. To see why condition (4.4) is necessary, consider the case $\frac{\alpha_{\pi}}{4b^{2}}<1$. Then by Lipschitz continuity of $g$, $\tau_{i+1}^{\ast}=|g(\tau_{i}^{\ast})-g(0)|\leq\frac{\alpha_{\pi}}{4b^{2}}|\tau_{i}^{\ast}-0|$ since $g(0)=0$ and $\tau_{i}^{\ast}>0$ for all $i$. By induction, we have that $\tau_{i}^{\ast}\leq\left(\frac{\alpha_{\pi}}{4b^{2}}\right)^{i}\tau_{0}^{\ast}$. Since $\sum_{i=1}^{\infty}\left(\frac{\alpha_{\pi}}{4b^{2}}\right)^{i}=\frac{1}{1-\frac{\alpha_{\pi}}{4b^{2}}}$ whenever $\frac{\alpha_{\pi}}{4b^{2}}<1$, it holds that $\sum_{i=1}^{\infty}\tau_{i}<\sum_{i=1}^{\infty}\tau_{i}^{\ast}\leq\frac{1}{1-\frac{\alpha_{\pi}}{4b^{2}}}$, so that it is impossible to construct a sequence of thresholds $\{\tau_{i}^{\ast}\}$ spanning an infinite time horizon. To see why condition (4.4) is sufficient for an infinite time control, we show that there exists a $\tau_{\infty}^{\ast}>0$ whenever $\tau_{1}\neq 0$. First note that $\frac{dg(0)}{d\tau}=\frac{\alpha_{\pi}}{4b^{2}}>0$, meaning $\tau=0$ is an unstable fixed point. Consider $h(\tau):=g(\tau)-\tau$ and note that $h(\tau)\rightarrow$-$\infty$ as $\tau\rightarrow\infty$ since $g(\tau)<\frac{1}{2b}\tan^{-1}\left(\frac{\alpha_{\pi}}{2b}\right)$ for all $\tau\geq 0$. Additionally, $h(0)=0$ and $\frac{dh(0)}{d\tau}=\frac{\alpha_{\pi}}{4b^{2}}-1>0$ under condition (4.4) and $\frac{dh(\tau)}{d\tau}\rightarrow-1$ as $\tau\rightarrow\infty$ monotonically. Then by continuity of $h$ and $g$, there must exist a $\tau>0$ such that $h(\tau)=0$, i.e. there exists a $\tau_{\infty}^{\ast}>0$. That is, we may construct a sequence of step sizes $\{\tau_{i}\}_{i=1,2,3,}$ with $0<\tau_{i}<\min(\tau_{1}^{\ast},\tau_{\infty}^{\ast}\neq 0)$ for all $i=1,2,3,\dots$ as a discretisation of the interval $[0,T]$ for all $T>0$, for which the recursion (3.15)-(3.16) converges to the continuous time ODE (3.18) as $\max_{i}\tau_{i}\rightarrow 0$.
For the case $\alpha_{h}-L_{\pi}>0$, following a similar line of reasoning yields
$\dot{c}_{t}=-2c_{t}^{2}-c_{t}+2b^{2}+\frac{\alpha_{\pi}}{2}.$
Since this case corresponds to a uniform in time preservation of log-concavity under the W flow, there are no further restrictions when considering the WFR flow.
Finally, we have that for both $\alpha_{h}-L_{\pi}<0$ and $\alpha_{h}-L_{\pi}>0$, we must solve
$\dot{c}_{t}$ $=-2c_{t}^{2}-c_{t}+r$ (3.21)
where $r>0$ in both cases. Using the change of variable $y_{t}=c_{t}+\frac{1}{4}$ and solving (3.21) via separation of variables, along with the fact that $\alpha_{t}=\frac{\alpha_{\pi}}{2}+c_{t}$ yields (4.5).
Appendix D Negativity of covariance term
D.1 Multivariate Gaussian
Lemma D.1.
Suppose $\nu(x)=\mathcal{N}(x;b,Q)$ and $\pi(x)=\mathcal{N}(x;m_{\pi},C_{\pi})$ with $0\prec Q\prec C_{\pi}$. Then
$\mathrm{Cov}_{\tilde{\nu}_{\tau}}\left(g(\nu),\left|\nabla g(\nu)\right|^{2}\right)<0$
where $g(\nu)=\log\frac{\nu(x)}{\pi(x)}$ and $\tilde{\nu}_{\tau}\propto\nu^{(1-\tau)}\pi^{\tau}$ for $\tau\in[0,1]$.
Proof.
Recall that
$\mathrm{Cov}_{\tilde{\nu}_{\tau}}\left(g(\nu),\left|\nabla g(\nu)\right|^{2}\right)=\mathbb{E}_{\tilde{\nu}_{\tau}}\left[\log\frac{\nu}{\pi}\left|\nabla\log\frac{\nu}{\pi}\right|^{2}\right]-\mathbb{E}_{\tilde{\nu}_{\tau}}\left[\log\frac{\nu}{\pi}\right]\mathbb{E}_{\tilde{\nu}_{\tau}}\left[\left|\nabla\log\frac{\nu}{\pi}\right|^{2}\right]$
Furthermore, $\tilde{\nu}_{\tau}=\mathcal{N}(x;b_{\tau},Q_{\tau})$ since $\tilde{\nu}_{\tau}$ corresponds to a unit time interpolation between two Gaussians. Starting with the second term in the above, and using the shorthand notation $B:=C_{\pi}^{-1}-Q^{-1}$, $a:=m_{\pi}-b_{\tau}$ and $l:=b-b_{\tau}$,
$\mathbb{E}_{\tilde{\nu}_{\tau}}\left[\log\frac{\nu}{\pi}\right]$ $=\tfrac{1}{2}\left(Tr[BQ_{\tau}]+c_{1}\right)$
$\mathbb{E}_{\tilde{\nu}_{\tau}}\left[\left|\nabla\log\frac{\nu}{\pi}\right|^{2}\right]$ $=Tr[B^{2}Q_{\tau}]+c_{2}$
where $c_{1}:=a^{\top}C_{\pi}^{-1}a-l^{\top}Q^{-1}l,c_{2}:=a^{\top}C_{\pi}^{-2}a+l^{\top}Q^{-2}l-2a^{\top}C_{\pi}^{-1}Q^{-1}l$ and using the identity that if $X\sim\mathcal{N}(m,C)$ then $\mathbb{E}[h(X)]=Tr[PC]+(m-b)^{\top}P(m-b)$ where $h(X):=(X-b)^{\top}P(X-b)$. Also, suppose $X\sim\tilde{\nu}_{\tau}$ and $Y=X-b_{\tau}\sim\mathcal{N}(0,Q_{\tau})$. Then,
$\mathbb{E}_{\tilde{\nu}}\left[\log\frac{\nu}{\pi}\left|\nabla\log\frac{\nu}{\pi}\right|^{2}\right]=\tfrac{1}{2}\left(\mathbb{E}[Y^{\top}BYY^{\top}B^{2}Y]+\mathbb{E}[Y^{\top}BYc_{2}]+4\mathbb{E}[Y^{\top}v_{1}Y^{\top}v_{2}]+\mathbb{E}[c_{1}Y^{\top}B^{2}Y]+c_{1}c_{2}\right)$
$=\tfrac{1}{2}\left(Tr[(2Q_{\tau}BQ_{\tau}+Q_{\tau}Tr[Q_{\tau}B])B^{2}]+c_{2}Tr[BQ_{\tau}]+4Tr[v_{1}v_{2}^{\top}Q_{\tau}]+c_{1}Tr[B^{2}Q_{\tau}]+c_{1}c_{2}\right)$
where $v_{1}:=C_{\pi}^{-1}a-Q^{-1}l,v_{2}:=C_{\pi}^{-2}a+Q^{-2}l-C_{\pi}^{-1}Q^{-1}l-Q^{-1}C_{\pi}^{-1}a$ and using that $Y$ has mean zero under $\tilde{\nu}_{\tau}$ and has zero third moment and using the identity $\mathbb{E}_{\nu}[Y^{\top}BYY^{\top}B^{2}Y]=Tr[(\Sigma(B+B^{\top})\Sigma+\Sigma Tr[\Sigma B])B^{2}]$ for any $Y\sim\mathcal{N}(0,\Sigma)$. Combining the above yields
$\mathrm{Cov}_{\tilde{\nu}_{\tau}}\left(g(\nu),\left|\nabla g(\nu)\right|^{2}\right)$ $=Tr[Q_{\tau}BQ_{\tau}B^{2}]+2Tr[v_{1}v_{2}^{\top}Q_{\tau}].$
As shown in [CHH+23, Lemma E.3], the covariance of the unit time interpolation between two Gaussians is given by
$Q_{\tau}$ $=((1-\tau)R^{-1}+\tau I)^{-1}C_{\pi}$
where $R:=QC_{\pi}^{-1}$ also then $B=C_{\pi}^{-1}(I-R^{-1})$. Therefore,
$Tr[Q_{\tau}BQ_{\tau}B^{2}]$ $=Tr[[(1-\tau)B^{-1}Q^{-1}+\tau B^{-1}C_{\pi}^{-1}]^{-2}B]$
$=Tr[(R-I)^{3}[(I+\tau(R-I))^{2}R]^{-1}C_{\pi}^{-1}]$
using the Woodbury identity $(R-I)^{-1}R=(I-R^{-1})^{-1}$. Recall that for $X,Y$ symmetric, with $X\prec 0,Y\succ 0$ it holds that $Tr[XY]<0$. As $f(R)$ is not guaranteed to be symmetric, we consider its symmetric part $\frac{1}{2}(f(R)+f(R)^{\top})$ and evaluate its eigenvalues. This holds since $Tr[XY]=\frac{1}{2}Tr[(X+X^{\top})Y]$ whenever $Y$ is symmetric. When $Q\prec C_{\pi}$, $\lambda_{i}\in(0,1)$ where $\lambda_{i}$ denotes the $i$th eigenvalue of $R$. Since $f(R)$ is a rational function in the space of matrices, we have that the $i$th eigenvalue of $f(R)$ is given by $f(\lambda_{i})$, i.e.
$f(\lambda_{i})=\frac{(\lambda_{i}-1)^{3}}{(1+\tau(\lambda_{i}-1))^{2}\lambda_{i}}<0,\quad\forall\enskip\tau\in[0,1]$
The same arguments hold for $f(R)^{\top}$ so that $\frac{1}{2}(f(R)+f(R)^{\top})$ is symmetric negative definite. Therefore, $Tr[f(R)C_{\pi}^{-1}]<0$.
Then for the second term, $Tr[v_{1}v_{2}^{\top}Q_{\tau}]$, let $S:=R^{\top}=C_{\pi}^{-1}Q$,
$Q_{\tau}=C_{\pi}(S^{-1}+\tau(I-S^{-1}))^{-1}$
and also
$v_{2}^{\top}Q_{\tau}$ $=(a^{\top}C_{\pi}^{-1}+l^{\top}Q^{-1}S^{-1}-l^{\top}Q^{-1}-a^{\top}C_{\pi}^{-1}S^{-1})(S^{-1}+\tau(I-S^{-1}))^{-1}$
$=(l^{\top}Q^{-1}-a^{\top}C_{\pi}^{-1})(S^{-1}-I)(S^{-1}+\tau(I-S^{-1}))^{-1}$
So finally,
$v_{1}v_{2}^{\top}Q_{\tau}$ $=(C_{\pi}^{-1}a-Q^{-1}l)(l^{\top}Q^{-1}-a^{\top}C_{\pi}^{-1})(S^{-1}-I)(S^{-1}+\tau(I-S^{-1}))^{-1}$
$=-v_{1}v_{1}^{\top}(S^{-1}-I)(S^{-1}+\tau(I-S^{-1}))^{-1}$
and by the same arguments as the first term, $v_{1}v_{1}^{\top}$ is symmetric positive definite and $(S^{-1}-I)(S^{-1}+\tau(I-S^{-1}))^{-1}\succ 0$ so that $Tr[v_{1}v_{2}^{\top}Q_{\tau}]<0$.
D.2 1D symmetric distributions
Lemma D.2.
Suppose $\nu,\pi\in\mathcal{P}()$ with $\nu(x)=\nu(2a-x)$ and $\pi(x)=\pi(2a-x)$. Let $g(\nu)=\log\frac{\nu(x)}{\pi(x)}$ be such that $\nabla^{2}g\preceq\alpha I$. Then
$\mathrm{Cov}_{\tilde{\nu}_{\tau}}\left(g(\nu),\left|\nabla g(\nu)\right|^{2}\right)<0$
where $\tilde{\nu}_{\tau}\propto\nu^{(1-\tau)}\pi^{\tau}$ for $\tau\in[0,1]$.
Proof.
It follows from the symmetry of $\nu,\pi$ that $\tilde{\nu}_{\tau}(x)\propto\nu(x)^{(1-\tau)}\pi(x)^{\tau}$ is also symmetric around $a$ and such that $\int_{x<a}\tilde{\nu}_{\tau}(x)dx=\int_{x>a}\tilde{\nu}_{\tau}(x)dx=1/2$ for all $\tau\in[0,1]$.
Let us denote h=gh=-g, so that $\text{Cov}_{\tilde{\nu}_{\tau}}(g,|\nabla g|^{2})=-\text{Cov}_{\tilde{\nu}_{\tau}}(h,|\nabla h|^{2})$. Under our assumptions, $h$ and $|\nabla h|^{2}$ are both decreasing on $\{x<a\}$ and increasing on $\{x>a\}$. Since $h$ and $|\nabla h|^{2}$ are both decreasing in $\{x<a\}$ and both increasing in $\{x>a\}$ we can apply Chebyshev’s integral inequality (e.g. [EGW09])
$\int_{\{x<a\}}h(x)|\nabla h(x)|^{2}\tilde{\nu}_{\tau}(x)dx\geq\frac{1}{\int_{\{x<a\}}\tilde{\nu}_{\tau}(x)dx}\left(\int_{\{x<a\}}h(x)\tilde{\nu}_{\tau}(x)dx\right)\left(\int_{\{x<a\}}|\nabla h(x)|^{2}\tilde{\nu}_{\tau}(x)dx\right)$
and likewise over $\{x>a\}$. Then,
$\text{Cov}_{\tilde{\nu}_{\tau}}(h,|\nabla h|^{2})$ $\geq\left(\frac{1}{\int_{\{x<a\}}\tilde{\nu}_{\tau}(x)dx}-2\right)\left(\int_{\{x<a\}}h(x)\tilde{\nu}_{\tau}(x)dx\right)\left(\int_{\{x<a\}}|\nabla h(x)|^{2}\tilde{\nu}_{\tau}(x)dx\right)$
$+\left(\frac{1}{\int_{\{x>a\}}\tilde{\nu}_{\tau}(x)dx}-2\right)\left(\int_{\{x>a\}}h(x)\tilde{\nu}_{\tau}(x)dx\right)\left(\int_{\{x>a\}}|\nabla h(x)|^{2}\tilde{\nu}_{\tau}(x)dx\right).$
where in the last line we used the symmetry assumption to obtain $\int_{\{x>a\}}\tilde{\nu}_{\tau}(x)h(x)dx=\int_{\{x<a\}}\tilde{\nu}_{\tau}(x)h(x)dx$ and similarly for $|\nabla h|^{2}$. Recalling that $\int_{x<a}\tilde{\nu}_{\tau}(x)dx=\int_{x>a}\tilde{\nu}_{\tau}(x)dx=1/2$ it follows that $\text{Cov}_{\tilde{\nu}_{\tau}}(g,|\nabla g|^{2})=\text{Cov}_{\tilde{\nu}_{\tau}}(h,|\nabla h|^{2})\geq 0$.